# Gaussian Grouping: Segment and Edit Anything in 3D Scenes

### v1

@@ 3.2.1 Anything Mask Input and Consistency
where $P$ contains the sampled Identity Encoding $e$ of a 3D Gaussian, while the set $Q=\{e^{\prime}_{1},e^{\prime}_{2},...,e^{\prime}_{k}\}$ consists of its $k$ nearest neighbors in 3D spatial space. We omit the softmax operation combined after linear layer $f$ for simplicity.
Training Objective Combined with the conventional 3D Gaussian Loss $\mathcal{L}_{\text{ori}}$ on image rendering [13], the total loss $\mathcal{L}_{\text{render}}$ for fully end-to-end training is
$\mathcal{L}_{\text{render}}=\mathcal{L}_{\text{ori}}+\mathcal{L}_{\text{id}}=\mathcal{L}_{\text{ori}}+\lambda_{\text{2d}}\mathcal{L}_{\text{2d}}+\lambda_{\text{3d}}\mathcal{L}_{\text{3d}}$ (4)
Figure 3: The grouped 3D Gaussians after training, where each group represents a specific instance / stuff of the 3D scene and can be fully decoupled. Our representation is efficient to support versatile downstream scene editing applications, where we design a Gaussian Operation List consists of simple operations like group deletion, group addition, finetuning Spherical Harmonic (SH) and exchanging 3D center locations.
3.3 Local Gaussian Editing
After the 3D Gaussian field training and grouping (Sec 3.2.2), as in Figure 3, we represent the whole 3D scene with a set of grouped 3D Gaussians. To perform various downstream local scene editing tasks, we propose the efficient Local Gaussian Editing. Thanks to the decoupled scene representation, instead of finetuing all 3D Gaussians, we freeze the properties for the most of the well-trained Gaussians and only adjust a small part of existing or newly added 3D Gaussians relevant to the editing targets. For 3D object removal, we simply delete the 3D Gaussians of the editing target. For 3D scene re-composition, we exchange the 3D location between two Gaussian groups. These two editing applications are direct with no parameter tuning.
For 3D object inpainting, we first delete the relevant 3D Gaussians and then add a small number of new Gaussians to be supervised by the 2D inpainting results by LAMA [41] during rendering. For 3D object colorization or style transfer, we only tune the color (SH) parameters of the corresponding Gaussian group, while fixing the 3D positions and sizes of them to preserve the learned 3D scene geometry.
The proposed local Gaussian editing scheme is time efficient, as shown in our experiment. Because of our fine-grained masks modeling, it also supports multiple concurrent local editings without interfering with each other or re-training the whole global 3D scene representation with new editing operation. Comparing to NeRF-based approaches [31, 17, 10], we show extensive visual comparisons for versatile scene editing cases in our experiment section and the Supp. file.
Figure 4: Ablation on the Identity Consistency across views, where we treat multi-view images as a video and associate the mask labels to generate a coherent segmentation labels [7] for training. We compare our rendering results (the 3rd row) to using cost-based linear assignment [39] during training (the 1st row), but found it leads to slow training and inferior testing performance in both reconstruction and segmentation.
@@ 4.1 Dataset and Experiment Setup
Figure 5: Segmentation comparison between LERF [14] and our Gaussian Grouping on the rendering view. The masks predicted by Gaussian Group contains much sharp and accurate boundary than LERF. Also, our approach is better at distinguish objects with similar colors, such as the “Green apple” prompt case. We input the corresponding text queries to Grounding DINO [24] to select the mask ID from our Anything Masks prediction. We adopt PCA to visualize the rendered Identity Encoding features in the rightmost column.
@@
Datasets To measure to segmentation or fine-grained localization accuracy in open-wold scene, we evolve the existing LERF-Localization [14] evaluation dataset and propose the LERF-Mask dataset, where we manually annotate three scenes from LERF-Localization with accurate masks instead of using coarse bounding boxes. For each 3D scene, we provide 7.7 text queries with corresponding GT mask label in average, more details are provided in the Supp. file.
@@
To evaluate the reconstruction quality, we tested our Gaussian Grouping on 7 of full 9 sets of scenes presented in Mip-Nerf360 [1], where the flowers and treehill are skipped due to the non-public access right. We also take diverse 3D scene cases from LLFF [28], Tanks&Temples [16] and Instruct-NeRF2NeRF [10] for visual comparison. We also adopt boundary metrics mBIoU for better segmentation quality measurement.
@@
Implementation Details We implement Gaussian Grouping based on Gaussian Splatting [13]. We add an 16 dimension identity encoding as a feature of each Gaussian and implement forward and backward cuda rasterization similar to rgb feature. The output classification linear layer has 16 input channels and 256 output channels. In training, $\lambda_{\text{2d}}=1.0$ and $\lambda_{\text{2d}}=2.0$. We use the Adam optimizer for both gaussians and linear layer, with a learning rate of 0.0025 for identity encoding and 0.0005 for linear layer. For 3D regularization loss, we choose $k=5$ and $m=1000$. All datasts are trained for 30K iterations on one A100 GPU.
For 3D object removal, we obtain the 3D Gaussians with classification of the selected instance id. Post-processing involves first removing the outliers and then using a convex hull to get all Gaussians corresponding to the object for deletion. For 3D object inpainting, we initialize Gaussians near the deletion region and finetune with the same LPIPS loss in inpainting mask as SPIn-NeRF [31].
4.2 Multiview Reconstruction and Segmentation
@@
Table 1: Comparison of Open Vocabulary Segmentation on LERF-Mask dataset. To compare fine-grained mask localization quality, we annotate the test views of three 3D scenes from LERF-Localization [14] dataset with accurate masks to replace its original coarse bounding boxes. We adopt the detected box from Grounding DINO [24] to select the mask ID in a 3D scene.
@@
Ablation on Mask Cross-view Association To study the effect of cross-view masks association [7] for input preparation, we replace the associated masks input to the individual masks predicted by SAM per image in Gaussian Grouping. We take the cost-based linear assignment strategy proposed in [39], and perform the visual comparison on rendering results in Figure 4. The linear assignment not only heavily slows down the whole training due to the assignment computation in each iteration, but also produces noised mask predictions. This is owing to the large gradients brought by the unstable cost-based liner assignment, especially at the initial training stage of the network, where masks prediction is nearly random. For 2K training iteration, linear assignment requires 1 hour but our associated mask input only requires 1 minute. Also, we compare the scene render reconstruction quality in Figure 4, where the appearance details of the rope attached to the ceiling (see the zoom region) is much better preserved by our method.
Table 2: Ablation of Identity Encoding on Mip-NeRF 360 [1] dataset with its seven public scenes. The joint training of the introduced Identity Encodings with original Gaussian parameters does not hurt the original Gaussians reconstruction quality.
@@
Gaussian Splatting [13] 28.69 0.870 0.182 $\sim$200
Gaussian Grouping $\checkmark$ 28.43 0.863 0.189 $\sim$170
@@
Figure 6: 3D Object removal on the Tanks&Temples dataset [16]. Comparing to DFFs [17], our Gaussian Grouping can remove the large-scale objects, such as tank, from the whole 3D scene with greatly reduced artifacts without leaving a blurry background.
Open-vocabulary Segmentation Comparison We compare the segmentation quality of Gaussian Grouping in 3D scenes with existing open-vocabulary 3D segmentation method, such as LERF [14]. In Table 1, the advantage of our Gaussian Grouping is significant, doubling the performance of LERF on both the “figurines” and “ramen” scenes. We also show the visual segmentation comparison in Figure 5, where the segmentation prediction by our method is much more accurate with clear boundary, while LERF-based similarity computing methods only provide a rough localization area. Since SAM does not support language prompt, we adopt the Grounding DINO [24] to identify the mask ID in a 2D image, and then pick the corresponding mask from our anything masks prediction (2nd column from the right).
Influence of the Identity Encoding In Table 2, we study the influence of our introduced Identity Encoding to original Gaussian Splatting’s 3D reconstruction performance and speed. The performance of Gaussian Grouping is on par with the original Gaussian Splatting method with negligible decrease, but it allows anything in the whole 3D scene to be segmented. By Identity Encoding, we lift the 2D zero-shot segmentation of SAM to 3D open world. Further, the grouped 3D Gaussians support a wide range of downstream editing tasks. We simply train all model components in an end-to-end manner jointly. This is different from Panoptic Lifting [20] which requires to block gradients from the segmentation branch back to the reconstruction branch.
Ablation on Identity Encoding Dimension We study the impact of dimension for our proposed Identity Encoding in Figure 4. To keep the training efficiency of Gaussian Grouping, we set the dimension of Identity Encoding to only 16 because it not only shows a good segmentation separation between objects but also keeps the training efficiency. Doubling dimension to 32 does not bring a better reconstruction quality compared to 16 but make training 1.3 times slower.
4.3 3D Object Removal
3D object removal is to completely delete an object from the 3D scene, where the background behind the removed instance / stuff can be noisy or have a hole because of no observation. In Figure 6, we compare the removal effect of our Gaussian Grouping with the Distilled Feature Fields (DFFs) [17]. For challenging scene cases with large objects, our method can clearly separate the 3D object from the background scene. While the performance of DFFs is limited by the quality of its CLIP-distilled features, which results in the complete foreground removal (Train case) or inaccurate region removal with obvious artifacts (Truck case).
@@
Figure 7: Comparison on 3D object inpainting cases, where SPIn-NeRF [31] requires 5h training while our method with better inpainting quality only needs 1 hour training and 20 minutes tuning.
Figure 8: Ablation on the Grouping Loss on the Bear inpainting case. The joint supervision of 2D and 3D losses addresses the “transparent bear issue”, which shows better Gaussian Grouping accuracy. After deleting the Gaussians belong to the bear, we detect the image hole region with no Gaussians projection covering and then use LAMA [41] to produce single view inpainting result to guide the learning for newly added inpainting Gaussians.
@@
Figure 9: Multi-object editing within the same image, where we concurrently perform 3D object removal for objects in the red and yellow circle, and colorization for the glove in the purple circle.
4.4 3D Object Inpainting
@@
Based on the 3D object removal, in 3D object inpainting task, we aim to further fill the “hole regions” due to missing of observation and make it a photo-realistic and view-consistent natural 3D scene. We first detect the regions which are invisible in all views after deletion, and then inpaint these “invisible regions” instead of the whole “2D object regions”. In each rendering view, we adopt the 2D inpainted image to guide the learning of newly introduced 3D Gaussian. In Figure 7, compared to SPIn-NeRF [31], the inpainting result of our Gaussian Grouping better preserves spatial detail and multi-view coherence.
Visual Ablation on the Grouping Losses In Figure 8, we study the effect of our grouping loss components, where solely using 2D Identity Loss will result in the ’transparent bear issue’. This is due to Gaussians inside bear will be occluded during rendering and cannot be supervised sufficiently. We address this issue by proposing 3D Identity Loss for joint training with the 2D loss.
4.5 3D Multi-Object Editing
@@
In Figure 9, we demonstrate multiple editing actions (like removing objects and colorization) on different Gaussian groups within the same 3D scene. These grouped 3D Gaussian representation enables simultaneous editing of several objects while maintaining non-interference among them.

### v2

@@ 3.1 Preliminaries: 3D Gaussian Splatting
where $P$ contains the sampled Identity Encoding $e$ of a 3D Gaussian, while the set $Q=\{e^{\prime}_{1},e^{\prime}_{2},...,e^{\prime}_{k}\}$ consists of its $k$ nearest neighbors in 3D Euclidean space. We omit the softmax operation combined after linear layer $f$ for simplicity.
Combined with the conventional 3D Gaussian Reconstruction Loss $\mathcal{L}_{\text{rec}}$ on image rendering [15], the total loss $\mathcal{L}_{\text{render}}$ for fully end-to-end training is
$\mathcal{L}_{\text{render}}=\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{id}}=\mathcal{L}_{\text{rec}}+\lambda_{\text{2d}}\mathcal{L}_{\text{2d}}+\lambda_{\text{3d}}\mathcal{L}_{\text{3d}}$ (4)
Figure 3: The grouped 3D Gaussians after training, where each group represents a specific instance / stuff of the 3D scene and can be fully decoupled. Our representation is efficient to support versatile downstream scene editing applications, where we design a Gaussian Operation List consisting of simple operations like group deletion, group addition, finetuning Spherical Harmonic (SH) and exchanging 3D center locations.
3.3 Gaussian Grouping for Scene Editing
After the 3D Gaussian field training and grouping (Sec 3.2), as in Figure 3, we represent the whole 3D scene with a set of grouped 3D Gaussians. To perform various downstream local scene editing tasks, we propose efficient Local Gaussian Editing. Thanks to the decoupled scene representation, instead of fine-tuning all 3D Gaussians, we freeze the properties for most of the well-trained Gaussians and only adjust a small part of existing or newly added 3D Gaussians relevant to the editing targets. For 3D object removal, we simply delete the 3D Gaussians of the editing target. For 3D scene re-composition, we exchange the 3D location between two Gaussian groups. These two editing applications are direct with no parameter tuning. For 3D object inpainting, we first delete the relevant 3D Gaussians and then add a small number of new Gaussians to be supervised by the 2D inpainting results by LaMa [46] during rendering. For 3D object colorization, we only tune the color (SH) parameters of the corresponding Gaussian group to preserve the learned 3D scene geometry. For 3D object style transfer, we further unfreeze the 3D positions and sizes to achieve more realistic results.
Our local Gaussian editing scheme is time-efficient, as shown in our experiment. Because of our fine-grained mask modeling, it also supports multiple concurrent local editings without interfering with each other or re-training the whole global 3D scene representation with new editing operations. Compared to NeRF-based approaches [34, 19, 12], we also provide extensive visual comparisons for scene editing cases and discussion on mask granularity in the Supp. file.
@@ 4.1 Dataset and Experiment Setup
4.1.1 Datasets
@@
To measure segmentation or fine-grained localization accuracy in open-world scene, we evolve the existing LERF-Localization [16] evaluation dataset and propose the LERF-Mask dataset, where we manually annotate three scenes from LERF-Localization with accurate masks instead of using coarse bounding boxes. For each 3D scene, we provide 7.7 text queries with corresponding GT mask labels on average. We also provide 3D panoptic segmentation results on Replica [45] and ScanNet [8] dataset. More details are in the Supp. file
@@
To evaluate the reconstruction quality, we tested our Gaussian Grouping on 7 of full 9 sets of scenes presented in Mip-NeRF 360 [1], where the flowers and treehill are skipped due to the non-public access right. We also take diverse 3D scene cases from LLFF [32], Tanks & Temples [18] and Instruct-NeRF2NeRF [12] for visual comparison on scene editing.
4.1.2 Implementation Details
@@
We implement Gaussian Grouping based on Gaussian Splatting [15]. We add a 16-dimension identity encoding to each Gaussian and implement forward and backward cuda rasterization similar to the RGB feature. The output classification linear layer has 16 input and 256 output channels. In training, $\lambda_{\text{2d}}=1.0$ and $\lambda_{\text{2d}}=2.0$. We use the Adam optimizer for both Gaussians and the linear layer, with a learning rate of 2.5e-3 for identity encoding and 5e-4 for the linear layer. For 3D regularization loss, we choose $k=5$ and $m=1000$. All datasets are trained for 30K iterations on one A100 GPU. Refer to our Supp. file for more details on scene editing and SAM’s mask granularity.
Figure 4: Ablation on the Identity Consistency across views, where we treat multi-view images as a video and associate the mask labels to generate coherent segmentation labels [7] for training. We founding using cost-based linear assignment [44] leads to slower training and inferior testing performance in both reconstruction and segmentation.
Figure 5: Robustness to input masks errors on Mip-NeRF 360 [1]. In the 2nd and 3rd columns (middle two views), SAM + DEVA fails to segment and associate the chair across frames. However, owing to the shared 3D Gaussian representation during reconstruction, Gaussian Grouping successfully corrects the error in mask labels and segments the black chair during the multi-view rendering.
4.2 Ablation Experiments
4.2.1 Ablation on Mask Cross-view Association
To study the effect of cross-view masks association [7] for input preparation, we replace the associated masks input to the individual masks predicted by SAM per image in Gaussian Grouping. We take the cost-based linear assignment strategy proposed in [44], and perform the visual comparison on rendering results in Figure 5. The linear assignment not only heavily slows down the whole training due to the assignment computation in each iteration, but also produces noisy mask predictions. This is owing to the large gradients brought by the unstable cost-based liner assignment, especially at the initial training stage of the network, where masks prediction is nearly random. For 2K training iteration, linear assignment requires 1 hour but our associated mask input only requires 1 minute. Also, we compare the scene render reconstruction quality in Figure 5, where the appearance details of the rope attached to the ceiling are much better preserved by our method.
4.2.2 Masks Association Errors Correction and Robustness
Gaussian Grouping can also correct the 2D segmentation errors produced by DEVA, as shown in Figure 5. Even taking the training input with partial 2D mask labels lost by tracking, Gaussian Grouping is robust in 3D association and can retrieve them based on the shared 3D Gaussian representation across different views.
Table 1: Influence of Identity Encoding on Mip-NeRF 360 [1] dataset with its 7 public scenes. The joint training of the introduced Identity Encodings does not hurt the original Gaussian reconstruction quality.
[…]
Baseline: Gaussian Splatting [15] - - 28.69 0.870 0.182 $\sim$200
Gaussian Grouping $\checkmark$ $\checkmark$ 28.43 0.863 0.189 $\sim$170
Table 2: Ablation of K of 3D Regularization Loss on the 3D object removal. RAcc: Object Removal Accuracy.
Model Gaussian Splatting Gaussian Grouping
K=0 K=1 k=2 K=5 K=10
PSNR 30.32 30.51 30.62 30.61 30.72 30.62
RAcc N/A 41.2% 40.5% 67.5% 76.6% 77.8%
4.2.3 Influence of the Identity Encoding
In Table 2, we study the influence of our introduced Identity Encoding on original Gaussian Splatting’s 3D reconstruction performance and speed. The performance of Gaussian Grouping is on par with the original Gaussian Splatting method with negligible decrease, but it allows anything in the whole 3D scene to be segmented. By Identity Encoding, we lift the 2D zero-shot segmentation of SAM to the 3D open world. Further, the grouped 3D Gaussians support a wide range of downstream editing tasks. We simply train all model components in an end-to-end manner jointly. This is different from Panoptic Lifting [22] which requires to block gradients from the segmentation branch back to the reconstruction branch.
4.2.4 Ablation on Identity Encoding Dimension
We study the impact of dimension for our Identity Encoding in Figure 5. To keep the training efficiency, we set the dimension of Identity Encoding to only 16 because it not only shows a good segmentation separation between objects but also keeps the training efficiency. Doubling the dimension to 32 does not bring a better reconstruction quality compared to 16 but make training 1.3 times slower.
Figure 6: Visual ablation of K in the 3D Regularization Loss on object removal editing of MipNeRF360. We remove Gaussians classified as lego with various K. We compute the convex hull of the removed 3D Gaussian points as the post process.
Figure 7: Ablation on the Grouping Loss on the Bear inpainting case. The joint supervision of 2D and 3D losses addresses the “transparent bear issue", which shows better Gaussian Grouping accuracy. After deleting the Gaussians belonging to the bear, we detect the image hole region with no Gaussians projection covering and then use LaMa [46] to produce a single view inpainting result to guide the learning for newly added inpainting Gaussians.
4.2.5 Ablation of 3D Regularization Loss
We perform ablation of K in our 3D Regularization Loss on the Kitchen dataset of Mip-NeRF 360 [1] to select K. As in Fig. 6 and Table 2, $K=5$ achieves both the best balance between the scene reconstruction and 3D object removal accuracy. Figure 6 further visualizes the influence of K during the 3D object removal process.
4.2.6 Visual Ablation on the Grouping Losses
In Figure 7, we study the effect of our grouping loss components, where solely using 2D Identity Loss will result in the ‘transparent bear issue’. This is due to Gaussians inside the bear being occluded during training and cannot be supervised sufficiently. We address it by proposing 3D Identity/Regularization Loss for joint training with the 2D loss.
4.3 3D Multi-view Segmentation
4.3.1 Open-vocabulary Segmentation Comparison
@@
We compare the segmentation quality of Gaussian Grouping in 3D scenes with the state-of-the-art open-vocabulary 3D segmentation methods, such as LERF [16] and SA3D [4]. To compare fine-grained mask localization quality, we annotate the test views of three 3D scenes from the LERF-Localization [16] dataset with accurate masks to replace its original coarse bounding boxes. In Table 4, the advantage of our Gaussian Grouping is significant, doubling the performance of LERF and SA3D on both the "figurines" and "ramen" scenes. We also show the visual segmentation comparison in Figure 8, where the segmentation prediction by our method is much more accurate with a clear boundary, while LERF-based similarity computing only provides a rough localization area. Since SAM does not support language prompts, both SA3D and our method adopt Grounding DINO [26] to identify the mask ID in a 2D image, and then pick the corresponding mask from our anything masks prediction.
4.3.2 3D Panoptic Segmentation Comparison
We compare the panoptic segmentation quality with Panoptic Lifting [44] in Table 4. We use the same semantic mask labels from Mask-DINO [23] with semantic information for both Panoptic Lifting [44] and our method. Gaussian Grouping outperforms Panoptic Lifting in both performance and speed.
Figure 8: Segmentation comparison between LERF [16] and our Gaussian Grouping on the rendering view. The masks predicted by Gaussian Grouping contain much sharper and more accurate boundaries than LERF. Also, our approach is better at distinguishing objects with similar colors, such as the “Green apple" prompt case.
4.4 3D Scene Editing
4.4.1 3D Object Removal
3D object removal is to completely delete an object from the 3D scene, where the background behind the removed instance / stuff can be noisy or have a hole because of no observation. In Figure 9, we compare the removal effect of our Gaussian Grouping with the Distilled Feature Fields (DFFs) [19]. For challenging scene cases with large objects, our method can clearly separate the 3D object from the background. While the performance of DFFs is limited by the quality of its CLIP-distilled features, which results in the complete foreground removal (Train case) or inaccurate region removal with obvious artifacts (Truck case).
Table 3: Comparison of Open Vocabulary Segmentation on LERF-Mask dataset. We adopt the detections from Grounding DINO [26] to select mask IDs in a 3D scene.
@@
Table 4: Panoptic segmentation comparison on novel views. For semantic information of the prediction, we adopt Mask-DINO [23] as the semantic mask label generator for each view. We refer to the reported setting from [9].
Model Replica ScanNet
@@
Panoptic Lifting [44] 66.22 64.34 $\sim$10 67.01 60.74 $\sim$10
Gaussian Grouping 71.15 66.52 $\sim$140 68.70 61.83 $\sim$150
@@
Figure 9: 3D Object removal on the Tanks & Temples dataset [18]. Compared to DFFs [19], our Gaussian Grouping can remove the large-scale objects, such as truck, from the 3D scene with greatly reduced artifacts w/o leaving a blurry background.
@@
Figure 10: Comparison on 3D object inpainting cases, where SPIn-NeRF [34] requires 5h training while our method with better inpainting quality only needs 1 hour training and 20 minutes finetuning.
@@
Figure 11: Multi-object editing within the same 3D scene, where we concurrently perform 3D object removal for objects in the red and green circles, and colorization for the glove in the purple circle.
Figure 12: Visual comparison of 3D object style transfer between Instruct- NeRF2NeRF [12] and Gaussian Grouping under the same instruction prompt “Turn the bear into a panda”.
4.4.2 3D Object Inpainting
@@
Based on the 3D object removal, in the 3D object inpainting task, we aim to further fill the “hole regions" due to missing observation across all views and make it a photo-realistic and view-consistent natural 3D scene. We first detect the regions which are invisible in all views after deletion, and then inpaint these “invisible regions” instead of the whole “2D object regions". In each rendering view, we adopt the 2D inpainted image to guide the learning of the newly introduced 3D Gaussian. In Figure 10, compared to SPIn-NeRF [34], the inpainting result of our Gaussian Grouping better preserves spatial detail and multi-view coherence.
4.4.3 3D Object Style Transfer
Gaussian Grouping supports 3D object style transfer efficiently. We compare with the recent Instruct-NeRF2NeRF [12] on 3D object style transfer in Figure 12, using the same instruction prompt “turn the bear into a panda" and the same image guidance by InstructPix2Pix [2]. Gaussian Grouping produces more coherent and natural transferred “bear” across views. Since our Gaussian Grouping models “anything masks” of the open-world 3D scene, including the bear, the spatial details of background regions (outside the bear) are also faithfully preserved by our method. While for Instruct-NeRF2NeRF, a large portion of background regions are unnecessarily getting blurry with degraded quality. We describe the detailed pipeline of 3D object style transfer in the Supp. file.
4.4.4 3D Multi-Object Editing
@@
In Figure 12, we demonstrate multiple editing actions (like removing objects and colorization) on different Gaussian groups of the 3D scene. This grouped 3D Gaussian representation enables concurrent editing of several objects while maintaining non-interference among them.
4.4.5 Limitation